A filament lamp operates at maximum brightness when connected to a 6.0 V supply. At maximum brightness, the current in the filament is 120 mA


A. (i) Calculate the resistance of the filament when it is operating at maximum brightness.


(ii) You have available a 24 V supply and a collection of resistors of a suitable power rating and with different values of resistance. Calculate the resistance of the resistor that is required to be connected in series with the supply such that the voltage across the filament lamp will be 6.0 V. 


1
Expert's answer
2021-12-28T13:36:45-0500


(i)

"V=IR"


"R=\frac{6}{120×10^{-3}}=50\Omega"


(ii)


Current in the circuit:

"I" "=\frac{Voltage\>across \>lamp}{Resistance \>of\>lamp}"


"=\frac{6}{50}=0.12A"



P.d across resistor "=24-6=18V"


Resistance of resistor "=\frac{18}{0.12}=150\Omega"